OpenStudy (anonymous):

Solve by using the addition method.. Provide the solution set: 3x^2-5y^2=-33 2x^2+5y^2=35

OpenStudy (cwrw238):

when you add these 2 equations the terms in y disappear to give 5x^2 = 2 solve for x (there will be 2 solutions) and plug these into one of the original equations to find the values of y.

OpenStudy (cwrw238):

can you finish this or need more help?

OpenStudy (anonymous):

But now I have another one that the y values do not disappear?!

OpenStudy (anonymous):

And I dont get them!!

OpenStudy (cwrw238):

ok post it

OpenStudy (anonymous):

2x^2-3y^2=-19 3x^2+2y^2-30

OpenStudy (cwrw238):

ok to do this you multiply equation 1 by 2 and equation 2 by 3 to give 4x^2 - 6y^2 = -38 9x^2 + 6y^2 = 90 is that = 30 in second equation?

OpenStudy (anonymous):

yes sorry lol

OpenStudy (cwrw238):

ok so if you add these 2 equations you will eliminate y

OpenStudy (cwrw238):

when you multiply equations be sure to multiply each term

OpenStudy (anonymous):

Thanks!!!!! I see it now!!!

OpenStudy (cwrw238):

good

OpenStudy (cwrw238):

that is what this website is for.

OpenStudy (anonymous):

So, would the answer be 2,3???